OpenStudy (anonymous):

IDENTIFY THE ERROR: 6p - (-2p - 5p)= -p

OpenStudy (anonymous):

more specific?

OpenStudy (anonymous):

you hsve to say why the -p is wrong. this has to do with factoring i think..

OpenStudy (anonymous):

6p-

OpenStudy (anonymous):

soo how is the 6p- wrong?

OpenStudy (anonymous):

6p + 2p + 5p = 13p ---> not -p

hero (hero):

Because everything on the left side would be changed to a plus

OpenStudy (anonymous):

6p-(-2p-5p)=6p-(-7p)=6p+7p=13p

OpenStudy (anonymous):

oohhh. i see now. thanks you (:

OpenStudy (anonymous):